HELP!!!!
Please compute
f(2k-7) if f(x)= 7x-40

Answer by stanbon(75887)

compute
f(2k-7) if f(x)= 7x-40
f(2k-7) = 7(2k-7)-40
=14k-49-40
=14k-89
---------
Cheers,
Stan H.